OpenStudy (anonymous):

what is the equation of the line that contains the points with (x, y) coordinates (-3, 7) and (5, -1)?

OpenStudy (anonymous):

Slope = m = y1 - y2 / x1 - x2 = 7 --1 / -3 -5 = 8 / -8 = -1 Slope point equation y - y1 = m(x - x1) y --1 = -1(x - 5) y + 1 = -x +5 y = -x + 4

OpenStudy (anonymous):

The equation is y = -x + 4 or x + y = 4 or x + y - 4 = 0.